### v1

@@ 1 Introduction
Table-top manipulation has rapidly advanced by learning from massive internet-scale datasets (Kim et al., 2024; Khazatsky et al., 2024; Brohan et al., 2023) and human demonstrations (Team, 2025; Qu et al., 2025). Yet, acquiring generalizable whole-body manipulation across embodiments, from dexterous hands (Yin et al., 2025b; Zhong et al., 2025; Qi et al., 2023) to humanoid whole-body control (Li et al., 2025b; Ze et al., 2025), remains prohibitively expensive due to hardware constraints, task complexity, and lack of large-scale embodiment-specific data (Zhang et al., 2025; Xu et al., 2025). On the other hand, abundant human motion data is readily available, including large-scale motion-capture datasets (Zhan et al., 2024; Mahmood et al., 2019), internet-scale video collections (Ren et al., 2025), and VR-based interfaces (Hoque et al., 2025). Recent advances in computer vision have made it possible to reconstruct 3D human body (Goel et al., 2023) and hand motion (Pavlakos et al., 2024), as well as object meshes (Xiang et al., 2025) and trajectories (Wen et al., 2024), directly from videos. These developments create a unique opportunity to leverage human motion as a scalable source of demonstrations for learning humanoid and dexterous robot control. However, a fundamental challenge arises: the embodiment gap - the mismatch in morphology, dynamics, and actuation between humans and robots - which leads to infeasible motion during transfer. This motivates our central research question:
@@
We formulate this as a physics-based retargeting problem (Reda et al., 2023): given kinematic human demonstrations, generate robot motions that (a) align the robot’s poses with human poses, (b) establish consistent contact with the environment, and (c) preserve the task objectives of the demonstrated behavior. The problem presents three key challenges: (a) Dynamical feasibility: There is a substantial embodiment gap between humans and robots, and reconstructed demonstrations from mesh estimation and state tracking are often noisy, making direct kinematic transfer infeasible. (b) Scalability and efficiency: The abundance of internet-scale human data requires approaches that are both computationally efficient and scalable to large datasets. (c) Robustness and missing contact information: Most human datasets lack the force and contact data required to ensure dynamical feasibility and preserve manipulation intent. Existing methods struggle to bridge these gaps: inverse-kinematics (IK) approaches (Qin et al., 2022; Yin et al., 2025a) are efficient but dynamically infeasible; reinforcement learning (RL) approaches (Li et al., 2025c; Lum et al., 2025) are general but requires expensive trajectory-specific training and tedious reward engineering; teleoperation (Yin et al., 2025b) is dynamically feasible but often labor-intensive and embodiment-specific.
@@ 2.2 Sampling for Physics-based Retargeting
Due to the contact-rich nature of physics-based retargeting, the optimization problem in equation˜1a is highly non-convex and often non-continuous. Sampling-based optimization (Mannor et al., 2003) provides a natural way to handle such landscapes, as it does not rely on smoothness or convexity assumptions. Intuitively, this resembles RL in that both rely on sampled trajectories from parallelized simulation to guide decision-making, but instead of updating a policy network, we directly use the samples to optimize the control sequence $U=u_{0:T-1}$. To this end, we adopt a sampling-based optimizer with an annealed sampling kernel (Xue et al., 2025; Pan et al., 2024):
@@
9:  Update $U^{i+1}$ using weighted average from equation˜3
@@
In contact-rich dexterous manipulation, the feasible solution set is typically narrow, so effective optimization requires a combination of coarse search to discover feasible contact modes and fine refinement to achieve stable contact. The annealed sampling covariance in equation˜3 implements this exploration-exploitation trade-off: $\beta_{1}$ controls the rate across outer iterations (global-to-local search over optimization updates), and $\beta_{2}$ controls the rate along the prediction horizon (allocating more or less perturbation across timesteps). As the schedule progresses, the effective sampling radius transitions from broad exploration to targeted exploitation around promising trajectories. Figure˜3 illustrates this effect: unlike standard sampling (e.g. MPPI) using a fixed search radius, annealed sampling shrinks the radius in later iterations, effectively reducing the variance of the sampled trajectories. To further improve speed, we adopt an early stopping strategy (Kamat et al., 2022), where optimization halts when improvements from further sampling become small. Algorithm˜1 presents a base sampling framework for physics-based retargeting.
@@ 2.3 Virtual Contact Guidance
This strategy connects to prior ideas: virtual object constraint (Mandi et al., 2025) $\|p_{k,t}^{\text{object}}-p_{k,t}^{\text{object,ref}}\|_{2}^{2}\leq\eta$, which expands feasibility around absolute object states in RL, and contact cost (Lakshmipathy et al., 2025), biasing optimization towards desired hand–object relative states. Our formulation reduces sampling complexity while preserving the intended contact sequence by maintaining relative hand–object relationships.
@@ 2.4 Trajectory Robustification
Handling imperfection in reference motion. To bridge the gap from reconstructed demonstrations to real hardware, we robustify trajectories against unknown or misspecified dynamics (e.g., friction, contact compliance) and reconstruction noise (e.g., object mesh estimation error, pose estimation error) that could otherwise make a nominal plan infeasible. Our approach optimizes a control sequence with a pessimistic (min–max) objective over a bounded parameter set $\mathcal{D}$, similar to domain randomization (DR) method (Tobin et al., 2017) but replacing expectation with worst-case cost to ensure universal feasibility:
@@
where $J(U,d)$ is the cost in equation˜1a under dynamics parameter $d$. In practice, $\mathcal{D}$ spans variations such as contact margin size, friction coefficients, and object mass. During optimization, each candidate sequence $U_{j}$ is rolled out over a mini-batch $d_{1:K}\sim\mathcal{D}$ and evaluated by $\max_{k\leq K}J(U_{j},d_{k})$, following robust sampling-based control (Williams et al., 2018). This formulation integrates seamlessly with the update rule in equation˜3, leverages GPU parallelization through batched rollouts, and has minimum computational overhead.
@@ 2.5 Physics-based Data Augmentation
Apart from generating feasible robot actions, another advantage of physics-based retargeting is the ability to systematically augment the retargeted data with diverse physics-aware behaviors. Starting from a single human demonstration, SPIDER is capable of diversifying a single behavior into a diverse set of physically feasible actions which could be used for downstream training (Mandlekar et al., 2023; Jiang et al., 2025; Yin et al., 2025b).
@@ 3.1 Experimental Setup
Dataset Selection. For dexterous manipulation tasks, we evaluate on 3 bimanual manipulation datasets: GigaHands (Fu et al., 2025), Oakink (Zhan et al., 2024) and ARCTIC (Li et al., 2023), comprising in total 1262 episodes and 2.4M frames across five distinct robotic hands and 103 different objects. For humanoid tasks, we choose commonly used LAFAN1 (Harvey et al., 2020), AMASS (Mahmood et al., 2019) for locomotion and OMOMO (Li et al., 2023) dataset for human-robot interaction. To the best of our knowledge, SPIDER is the first universal retargeting method that can handle both dexterous manipulation and humanoid whole-body control tasks at this scale.
@@
Figure 5: Specifications of robots used in evaluation. SPIDER supports both dexterous hand and humanoid robot. The significant variations in DoF, dimensions, and finger count demonstrate the cross-embodiment generalizability of our approach. We employ a simulated 12-DoF configuration (Li et al., 2025c) of the Inspire and Ability hands, removing the joint constraints compared to their real-world versions.
Evaluation Metrics. For dexterous manipulation and humanoid-object interaction tasks, we evaluate tracking performance on the object motion, since the goal is to move the object rather than to precisely track the robot hand. For object tracking error, we compute per-step averaged rotation error $E_{\text{rot}}=\frac{1}{T}\sum_{t=1}^{T}\arccos\left(2\langle q_{{\text{obj}},t},q_{{\text{obj}},t}^{\text{ref}}\rangle^{2}-1\right)$ and position error $E_{\text{pos}}=\frac{1}{T}\sum_{t=1}^{T}\|p_{{\text{obj}},t}-p_{{\text{obj}},t}^{\text{ref}}\|_{2}$, where $T$ is the number of timesteps, $q_{{\text{obj}},t}$ and $p_{{\text{obj}},t}$ are the object quaternion and position at timestep $t$, respectively. We follow the same evaluation setting as ManipTrans (Li et al., 2025c) for fair comparison, where success is defined using: (1) the rotation error $E_{\text{rot}}<0.5$ rad, and (2) the per-step translation error $E_{\text{pos}}<3$ cm. For humanoid locomotion, the joint tracking error, pelvis position error and pelvis orientation error are computed.
@@ 3.2 Ablation Study
To ensure representative evaluation while maintaining computational feasibility, we conduct our ablation study on a collection of challenging manipulation tasks from established datasets: (1) for Oakink (Zhan et al., 2024), we evaluate on 8 distinct bimanual two-object manipulation tasks that were previously benchmarked in ManipTrans (Li et al., 2025c); and (2) for GigaHands (Fu et al., 2025), we utilize their released example dataset comprising five diverse manipulation trajectories.
Baseline Selection. We compare four approaches: (1) Kinematic retargeting: fingertip inverse kinematics as an initial-guess quality baseline; (2) vanilla Sampling (based on MPPI (Howell et al., 2022)): standard sampling-based control; (3) Sampling with an annealed kernel (similar to DIAL-MPC (Xue et al., 2025)): differs from our method only in the absence of virtual guidance; and (4) Sampling with an annealed kernel and virtual contact guidance (our full method). We exclude RL-based comparisons here since they produce policies rather than direct control sequences, making direct comparison inappropriate. RL methods are evaluated separately in section˜3.3 for motion quality assessment.
@@ 3.3 Dexterous Manipulation Retargeting Results
Oakink (Zhan et al., 2024) $1022$ $0.413$ $0.459$ $\bm{0.479}$ $0.431$ $0.422$
GigaHands (Fu et al., 2025) $756$ $0.741$ $0.810$ $\bm{0.879}$ $0.706$ $0.812$
@@
Baselines and Reference Methods. We compare against state-of-the-art policy-learning systems where available: ManipTrans (Li et al., 2025c) on Oakink and DexMachina (Mandi et al., 2025) on ARCTIC. Unlike these methods, which train policies, our approach directly optimizes control, so the comparison serves only as a performance reference. Generation efficiency is measured by Frames Per Second (FPS), defined as the number of frames in a trajectory divided by computation time. Matching FPS to the reference trajectory frequency indicates real-time generation; lower values indicate slower speeds. Section˜3.3 show that SPIDER achieves competitive success rates on simpler Oakink tasks but lags on more complex ARCTIC manipulations. Its main advantage is speed: trajectory generation is an order of magnitude faster, enabling large-scale dataset retargeting and potential online applications.
@@ 3.4 Humanoid Whole-Body Control Retargeting Results
As a universal physics-based retargeting method, SPIDER is versatile and can be applied to various robot embodiments, including humanoid robots. To adapt SPIDER for humanoid robots, virtual contact guidance is implemented between the robot’s feet and the floor. In figure˜8, we showcase the application of SPIDER on the humanoid AMASS (Mahmood et al., 2019) and LAFAN1 (Harvey et al., 2020) datasets. The retargeting process corrects artifacts such as foot penetration and slipping, enabling the robot to execute highly dynamic motions, as illustrated in figure˜8, which shows an open-loop rollout of the retargeted control and highlights the dynamical feasibility of the retargeted motion. We compare SPIDER with popular kinematics-based retargeting method, named general motion retargeting (GMR) (Araujo et al., 2025).
@@
Table 4: Tracking Error and FPS Comparison on Humanoid Datasets. Joint Err.: mean joint angle difference (degrees). Pos. Err.: mean end-effector position error (cm). Ori. Err.: mean end-effector orientation error (degrees). Obj. Err.: mean object pose error (cm or degrees). FPS: trajectory generation speed. Our method achieves the lowest tracking errors and highest FPS compared to GMR and OmniRetarget across all datasets.
@@ 4.1 Retargeting from Single RGB Camera
To demonstrate the flexibility of SPIDER for upstreaming data, we evaluate it on converting single RGB camera video into executable robot trajectories. The pipeline consists of: (1) 3D mesh reconstruction with Trellis (Xiang et al., 2025), (2) hand pose estimation with HAMER (Pavlakos et al., 2024) to obtain MANO parameters, and (3) object pose tracking with FoundationPose (Wen et al., 2024) for 6D trajectories. Due to the occlusion and limited resolution, the human and object motion are more noisy than the dataset, which requires more robust physics grounding. As shown in figure˜9, we validate on single-hand manipulation tasks like pouring with a cup and installing a bolt. Real-world noise, artifacts, and penetrations are corrected through physics grounding in SPIDER, and the retargeted trajectories are directly executable on physical robots.

### v2

@@ 1 Introduction
Table-top manipulation has rapidly advanced by learning from massive internet-scale datasets (kimOpenVLAOpenSourceVisionLanguageAction2024; khazatsky2024droid; brohan2023rt2visionlanguageactionmodelstransfer) and human demonstrations (rdt2; qu2025eo1interleavedvisiontextactionpretraining). Yet, acquiring generalizable whole-body manipulation across embodiments, from dexterous hands (yinDexterityGenFoundationController2025; zhongDexGraspVLAVisionLanguageActionFramework2025; qi2023general) to humanoid whole-body control (liAMOAdaptiveMotion2025; zeTWISTTeleoperatedWholeBody2025), remains prohibitively expensive due to hardware constraints, task complexity, and lack of large-scale embodiment-specific data (zhangDOGloveDexterousManipulation2025; xuDexUMIUsingHuman2025). On the other hand, abundant human motion data is readily available, including large-scale motion-capture datasets (zhanOAKINK2DatasetBimanual2024; mahmoodAMASSArchiveMotion2019), internet-scale video collections (renMotionTracksUnified2025), and VR-based interfaces (hoqueEgoDexLearningDexterous2025). Recent advances in computer vision have made it possible to reconstruct 3D human body (goelHumans4DReconstructing2023) and hand motion (pavlakosReconstructingHands3D2023), as well as object meshes (xiangStructured3DLatents2025) and trajectories (wenFoundationPoseUnified6D2024), directly from videos. These developments create a unique opportunity to leverage human motion as a scalable source of demonstrations for learning humanoid and dexterous robot control. However, a fundamental challenge arises: the embodiment gap - the mismatch in morphology, dynamics, and actuation between humans and robots - which leads to infeasible motion during transfer. This motivates our central research question:
@@
We formulate this as a physics-based retargeting problem (redaPhysicsbasedMotionRetargeting2023): given kinematic human demonstrations, generate robot motions that (a) align the robot’s poses with human poses, (b) establish consistent contact with the environment, and (c) preserve the task objectives of the demonstrated behavior. The problem presents three key challenges: (a) Dynamical feasibility: There is a substantial embodiment gap between humans and robots, and reconstructed demonstrations from mesh estimation and state tracking are often noisy, making direct kinematic transfer infeasible. (b) Scalability and efficiency: The abundance of internet-scale human data requires approaches that are both computationally efficient and scalable to large datasets. (c) Robustness and missing contact information: Most human datasets lack the force and contact data required to ensure dynamical feasibility and preserve manipulation intent. Existing methods struggle to bridge these gaps: inverse-kinematics (IK) approaches (qinOneHandMultiple2023; yinGeometricRetargetingPrincipled2025) are efficient but dynamically infeasible; reinforcement learning (RL) approaches (liManipTransEfficientDexterous2025; lumCrossingHumanRobotEmbodiment2025) are general but requires expensive trajectory-specific training and tedious reward engineering; teleoperation (yinDexterityGenFoundationController2025) is dynamically feasible but often labor-intensive and embodiment-specific.
@@ 2.2 Sampling for Physics-based Retargeting
Due to the contact-rich nature of physics-based retargeting, the optimization problem in equation˜1a is highly non-convex and often non-continuous. Sampling-based optimization (mannorCrossEntropyMethod) provides a natural way to handle such landscapes, as it does not rely on smoothness or convexity assumptions. Intuitively, this resembles RL in that both rely on sampled trajectories from parallelized simulation to guide decision-making, but instead of updating a policy network, we directly use the samples to optimize the control sequence $U=u_{0:T-1}$. To this end, we adopt a sampling-based optimizer with an annealed sampling kernel (xueFullOrderSamplingBasedMPC2024; panModelBasedDiffusionTrajectory2024):
@@
9:  Update $U^{i+1}$ using weighted average from equation˜2
@@
In contact-rich dexterous manipulation, the feasible solution set is typically narrow, so effective optimization requires a combination of coarse search to discover feasible contact modes and fine refinement to achieve stable contact. The annealed sampling covariance in equation˜2 implements this exploration-exploitation trade-off: $\beta_{1}$ controls the rate across outer iterations (global-to-local search over optimization updates), and $\beta_{2}$ controls the rate along the prediction horizon (allocating more or less perturbation across timesteps). As the schedule progresses, the effective sampling radius transitions from broad exploration to targeted exploitation around promising trajectories. Figure˜3 illustrates this effect: unlike standard sampling (e.g. MPPI) using a fixed search radius, annealed sampling shrinks the radius in later iterations, effectively reducing the variance of the sampled trajectories. To further improve speed, we adopt an early stopping strategy (kamatBITKOMOCombiningSampling2022), where optimization halts when improvements from further sampling become small. Algorithm˜1 presents a base sampling framework for physics-based retargeting.
@@ 2.3 Virtual Contact Guidance
This strategy connects to prior ideas: virtual object constraint (mandiDexMachinaFunctionalRetargeting2025) $\|p_{k,t}^{\text{object}}-p_{k,t}^{\text{object,ref}}\|_{2}^{2}\leq\eta$, which expands feasibility around absolute object states in RL, and contact cost (lakshmipathyKinematicMotionRetargeting2024), biasing optimization towards desired hand–object relative states. Our formulation reduces sampling complexity while preserving the intended contact sequence by maintaining relative hand–object relationships.
@@ 2.4 Trajectory Robustification
Handling imperfection in reference motion. To bridge the gap from reconstructed demonstrations to real hardware, we robustify trajectories against unknown or misspecified dynamics (e.g., friction, contact compliance) and reconstruction noise (e.g., object mesh estimation error, pose estimation error) that could otherwise make a nominal plan infeasible. Our approach optimizes a control sequence with a pessimistic (min–max) objective over a bounded parameter set $\mathcal{D}$, similar to domain randomization (DR) method (tobinDomainRandomizationTransferring2017a) but replacing expectation with worst-case cost to ensure universal feasibility:
@@
where $J(U,d)$ is the cost in equation˜1a under dynamics parameter $d$. In practice, $\mathcal{D}$ spans variations such as contact margin size, friction coefficients, and object mass. During optimization, each candidate sequence $U_{j}$ is rolled out over a mini-batch $d_{1:K}\sim\mathcal{D}$ and evaluated by $\max_{k\leq K}J(U_{j},d_{k})$, following robust sampling-based control (williamsRobustSamplingBased2018). This formulation integrates seamlessly with the update rule in equation˜2, leverages GPU parallelization through batched rollouts, and has minimum computational overhead.
@@ 2.5 Physics-based Data Augmentation
Apart from generating feasible robot actions, another advantage of physics-based retargeting is the ability to systematically augment the retargeted data with diverse physics-aware behaviors. Starting from a single human demonstration, SPIDER is capable of diversifying a single behavior into a diverse set of physically feasible actions which could be used for downstream training (mandlekar2023mimicgen; jiangDexMimicGenAutomatedData2025; yinDexterityGenFoundationController2025).
@@ 3.1 Experimental Setup
Dataset Selection. For dexterous manipulation tasks, we evaluate on 3 bimanual manipulation datasets: GigaHands (fuGigaHandsMassiveAnnotated2024), Oakink (zhanOAKINK2DatasetBimanual2024) and ARCTIC (li2023object), comprising in total 1262 episodes and 2.4M frames across five distinct robotic hands and 103 different objects. For humanoid tasks, we choose commonly used LAFAN1 (harvey2020robust), AMASS (mahmoodAMASSArchiveMotion2019) for locomotion and OMOMO (li2023object) dataset for human-robot interaction. To the best of our knowledge, SPIDER is the first universal retargeting method that can handle both dexterous manipulation and humanoid whole-body control tasks at this scale.
@@
Figure 5: Specifications of robots used in evaluation. SPIDER supports both dexterous hand and humanoid robot. The significant variations in DoF, dimensions, and finger count demonstrate the cross-embodiment generalizability of our approach. We employ a simulated 12-DoF configuration (liManipTransEfficientDexterous2025) of the Inspire and Ability hands, removing the joint constraints compared to their real-world versions.
Evaluation Metrics. For dexterous manipulation and humanoid-object interaction tasks, we evaluate tracking performance on the object motion, since the goal is to move the object rather than to precisely track the robot hand. For object tracking error, we compute per-step averaged rotation error $E_{\text{rot}}=\frac{1}{T}\sum_{t=1}^{T}\arccos\left(2\langle q_{{\text{obj}},t},q_{{\text{obj}},t}^{\text{ref}}\rangle^{2}-1\right)$ and position error $E_{\text{pos}}=\frac{1}{T}\sum_{t=1}^{T}\|p_{{\text{obj}},t}-p_{{\text{obj}},t}^{\text{ref}}\|_{2}$, where $T$ is the number of timesteps, $q_{{\text{obj}},t}$ and $p_{{\text{obj}},t}$ are the object quaternion and position at timestep $t$, respectively. We follow the same evaluation setting as ManipTrans (liManipTransEfficientDexterous2025) and DexMachina (mandiDexMachinaFunctionalRetargeting2025), where we report task success only concerning object motion. Task success is defined using: (1) the rotation error $E_{\text{rot}}<0.5$ rad, and (2) the average per-step translation error $E_{\text{pos}}<0.1$ m. For humanoid locomotion, the joint tracking error, pelvis position error and pelvis orientation error are computed.
@@ 3.2 Ablation Study
To ensure representative evaluation while maintaining computational feasibility, we conduct our ablation study on a collection of challenging manipulation tasks from established datasets: (1) for Oakink (zhanOAKINK2DatasetBimanual2024), we evaluate on 7 distinct bimanual two-object manipulation tasks that were previously benchmarked in ManipTrans (liManipTransEfficientDexterous2025); and (2) for GigaHands (fuGigaHandsMassiveAnnotated2024), we utilize their released example dataset comprising 3 manipulation trajectories.
Baseline Selection. We compare four approaches: (1) Kinematic retargeting: fingertip inverse kinematics as an initial-guess quality baseline; (2) vanilla Sampling (based on MPPI (howellPredictiveSamplingRealtime2022)): standard sampling-based control; (3) Sampling with an annealed kernel (similar to DIAL-MPC (xueFullOrderSamplingBasedMPC2024)): differs from our method only in the absence of virtual guidance; and (4) Sampling with an annealed kernel and virtual contact guidance (our full method). We exclude RL-based comparisons here since they produce policies rather than direct control sequences, making direct comparison inappropriate. RL methods are evaluated separately in section˜3.3 for motion quality assessment.
@@ 3.3 Dexterous Manipulation Retargeting Results
Oakink (zhanOAKINK2DatasetBimanual2024) $1022$ $0.413$ $0.459$ $\bm{0.479}$ $0.431$ $0.422$
GigaHands (fuGigaHandsMassiveAnnotated2024) $756$ $0.741$ $0.810$ $\bm{0.879}$ $0.706$ $0.812$
@@
Baselines and Reference Methods. We compare against state-of-the-art policy-learning systems where available: ManipTrans (liManipTransEfficientDexterous2025) on Oakink and DexMachina (mandiDexMachinaFunctionalRetargeting2025) on ARCTIC. Unlike these methods, which train policies, our approach directly optimizes control, so the comparison serves only as a performance reference. Generation efficiency is measured by Frames Per Second (FPS), defined as the number of frames in a trajectory divided by computation time. Matching FPS to the reference trajectory frequency indicates real-time generation; lower values indicate slower speeds. Section˜3.3 show that SPIDER achieves competitive success rates on simpler Oakink tasks but lags on more complex ARCTIC manipulations. Its main advantage is speed: trajectory generation is an order of magnitude faster, enabling large-scale dataset retargeting and potential online applications.
@@ 3.4 Humanoid Whole-Body Control Retargeting Results
As a universal physics-based retargeting method, SPIDER is versatile and can be applied to various robot embodiments, including humanoid robots. To adapt SPIDER for humanoid robots, virtual contact guidance is implemented between the robot’s feet and the floor. In figure˜8, we showcase the application of SPIDER on the humanoid AMASS (mahmoodAMASSArchiveMotion2019) and LAFAN1 (harvey2020robust) datasets. The retargeting process corrects artifacts such as foot penetration and slipping, enabling the robot to execute highly dynamic motions, as illustrated in figure˜8, which shows an open-loop rollout of the retargeted control and highlights the dynamical feasibility of the retargeted motion. We compare SPIDER with popular kinematics-based retargeting method, named general motion retargeting (GMR) (araujo2025retargetingmattersgeneralmotion).
@@
Table 4: Tracking Error and FPS Comparison on Humanoid Datasets. Joint Err.: mean joint angle difference (degrees). Pos. Err.: mean end-effector position error (cm). Ori. Err.: mean end-effector orientation error (degrees). Obj. Err.: mean object pose error (cm or degrees). FPS: trajectory generation speed. Our method achieves the lowest tracking errors and highest FPS compared to GMR across all datasets.
@@ 4.1 Retargeting from Single RGB Camera
To demonstrate the flexibility of SPIDER for upstreaming data, we evaluate it on converting single RGB camera video into executable robot trajectories. The pipeline consists of: (1) 3D mesh reconstruction with Trellis (xiangStructured3DLatents2025), (2) hand pose estimation with HAMER (pavlakosReconstructingHands3D2023) to obtain MANO parameters, and (3) object pose tracking with FoundationPose (wenFoundationPoseUnified6D2024) for 6D trajectories. Due to the occlusion and limited resolution, the human and object motion are more noisy than the dataset, which requires more robust physics grounding. As shown in figure˜9, we validate on single-hand manipulation tasks like pouring with a cup and installing a bolt. Real-world noise, artifacts, and penetrations are corrected through physics grounding in SPIDER, and the retargeted trajectories are directly executable on physical robots.